OpenStudy (anonymous):

y varies inversely as x, and y=50 when x=10. What is the value of y when x=20

OpenStudy (gorv):

this means y=a/x where a is constant y=50 and x=10 50=a/10 a=500 y=500/x at x=20 y=500/20 y=25

OpenStudy (gorv):

@tyianawooten